OpenStudy (anonymous):

A soccer ball is given a kick – it ends up going with an initial velocity of 28 m/s at an angle of 35 above the horizontal. Neglecting air friction, (a) how long was the ball in the air, and (b) what is the horizontal displacement of the kick?

OpenStudy (anonymous):

Are you familiar with components? If you do, \(S_y=u_y t +\frac{1}{2} a_y t^2\) as \(S_y=0\), i.e. it goes up and then back down to the ground, \(u_y=28 \sin 35^o ms^{-1}\), \(a_y =-9.8\) plug it all in to get t. \(S_x=u_x t +\frac{1}{2} a_x t^2\) where \(a_x =0\), t is found from the first, and \(u_x =28 \cos 35^o ms^{-1}\) plug it all in to get \(S_x\) do you need further help?